# Harnessing the Universal Geometry of Embeddings

### v2

@@ Abstract
The ability to translate unknown embeddings into a different space while preserving their geometry has serious implications for the security of vector databases. An adversary with access only to embedding vectors can extract sensitive information about the underlying documents, sufficient for classification and attribute inference.
@@ 1 Introduction
The Platonic Representation Hypothesis [17] conjectures that all image models of sufficient size have the same latent representation. We propose a stronger, constructive version of this hypothesis for text models: the universal latent structure of text representations can be learned and, furthermore, harnessed to translate representations from one space to another without any paired data or encoders.
@@
vec2vec is the first method to successfully translate embeddings from the space of one model to another without any paired data. vec2vec translations achieve cosine similarity as high as $0.92$ to the ground-truth vectors in their target embedding spaces and perfect matching on over $8000$ shuffled embeddings (without access to the set of possible matches in advance).
To show that our translations preserve not only the relative geometry of embeddings but also the semantics of underlying inputs, we extract information from them using zero-shot attribute inference and inversion, without any knowledge of the model that produced the original embeddings!
@@ 2 Problem formulation: unsupervised embedding translation
Figure 3: Unsupervised embedding translation. With access to only $u_{i}=M_{1}(d_{i})$, vec2vec seeks to generate a translation $F(u_{i})$ that is close in $M_{2}$’s embedding space to the ideal embedding $v_{i}=M_{2}(d_{i})$ without access to $d_{i}$, $v_{i}$, $M_{1}$, or $M_{2}$.
@@
Correspondence methods cannot be used. There is significant prior research on the problem of matching or correspondence between sets of embedding vectors [1, 43, 7, 48]. None of these approaches are applicable in our setting because they assume the existence of two (or more) sets of embedding vectors produced by different encoders from the same inputs. In other words, for each unknown vector, there must already exist a set of candidate vectors in a different embedding, one of which is the correct match. In practice, it is unrealistic to expect that such a database be available.  [^1] Our code is available on GitHub.
In our setting, where we do not assume access to the encoder $M_{1}$, nor any additional representations of the unknown documents $\{d_{1},\ldots,d_{n}\}$ other than their embeddings $u_{i}=M_{1}(d_{i})$, the problem is strictly harder and requires unsupervised translation from $M_{1}$ to $M_{2}$. Any success at unsupervised translation will come solely from the common geometric structure of $M_{1}$’s and $M_{2}$’s output spaces.
@@ 4.1 Preliminaries
Training. Unless otherwise specified, each vec2vec is trained on two sets of embeddings generated from disjoint sets of $1$ million $64$-token sequences sampled from NQ (see Appendix˜C for experiments with fewer embeddings). Due to GAN instability [47], we select the best of multiple initializations and leave more robust training to future work.
@@ 4.2 Evaluating translation
Let $u_{i}=M_{1}(d_{i})$ and $v_{i}=M_{2}(d_{i})$ denote the source and target embeddings of the same input $d_{i}$. The goal of translation is to generate a vector that is as close to $v_{i}$ is possible. We say that $(u_{i},v_{j})$ are “aligned” by the translator $F$ if $v_{j}$ is the closest embedding to $F(u_{i})$: $j=\arg\min_{k}\cos\bigl(F(u_{i}),v_{k}\bigr).$ A perfect translator $F^{*}$ satisfies $i=\arg\min_{k}\cos\bigl(F^{*}(u_{i}),v_{k}\bigr)$ for all $i$.
@@
vec2vec is the first unsupervised embedding translator, thus there is no direct baseline. As our Naïve baseline, we simply use $F(x)=x$ to measure geometric similarity between embedding spaces. The second (pseudo)baseline is Oracle-aided optimal assignment. It assumes that candidate targets are known and is thus strictly easier than vec2vec and the Naïve baseline. We solve optimal assignment, $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{n}\cos(u_{i},v_{\pi(i)})$, via either the Hungarian, Earth Mover’s Distance, Sinkhorn, or Gromov-Wasserstein algorithms, choosing the best-performing solver per experiment.
@@ 4.3 Evaluating information extraction
The second method is embedding inversion that recovers text inputs from embeddings. Since [37] requires a pre-trained inversion model for each embedding space, we use [59] instead to generate an approximation $d^{\prime}_{i}$ of $d_{i}$ from $F(M_{1}(d_{i}))$ in a zero-shot manner. We measure the extracted information using LLM judge accuracy: the fraction of translated embeddings for which GPT-4o determines that $d^{\prime}$ reveals information in $d$. See Appendix˜D for our prompt.
@@
vec2vec Naïve Baseline OA Baseline
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ Top-1 $\uparrow$ Rank $\downarrow$
gran. gtr 0.80 (0.0) 0.99 1.19 (0.1) -0.03 (0.0) 0.00 4168.73 (9.2) 0.00 4096.75 (9.2)
gte 0.87 (0.0) 0.95 1.18 (0.0) 0.01 (0.0) 0.00 4088.58 (9.2) 0.00 4069.91 (9.2)
stel. 0.79 (0.0) 0.98 1.05 (0.0) 0.01 (0.0) 0.00 4208.26 (9.2) 0.00 4096.75 (9.2)
e5 0.85 (0.0) 0.98 1.11 (0.0) 0.02 (0.0) 0.00 4111.60 (9.2) 0.00 4096.75 (9.2)
gtr gran. 0.81 (0.0) 0.99 1.02 (0.0) -0.03 (0.0) 0.00 4169.76 (9.2) 0.00 4096.66 (9.2)
gte 0.87 (0.0) 0.93 2.31 (0.1) 0.04 (0.0) 0.00 4080.92 (9.2) 0.00 4079.92 (9.2)
stel. 0.80 (0.0) 0.99 1.03 (0.0) 0.00 (0.0) 0.00 4198.78 (9.2) 0.00 4096.69 (9.2)
e5 0.83 (0.0) 0.84 2.88 (0.2) 0.03 (0.0) 0.00 4082.84 (9.2) 0.00 4066.42 (9.2)
gte gran. 0.75 (0.0) 0.95 1.22 (0.0) 0.01 (0.0) 0.00 4079.81 (9.3) 0.00 4069.23 (9.2)
gtr 0.75 (0.0) 0.91 2.64 (0.1) 0.04 (0.0) 0.00 4084.15 (9.2) 0.00 4078.45 (9.2)
stel. 0.89 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.87 (0.0) 0.99 5.19 (0.5) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
stel. gran. 0.80 (0.0) 0.98 1.08 (0.0) 0.01 (0.0) 0.00 4209.08 (9.3) 0.00 4096.88 (9.2)
gtr 0.82 (0.0) 1.00 1.10 (0.0) 0.00 (0.0) 0.00 4192.31 (9.2) 0.00 4096.63 (9.2)
gte 0.92 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.86 (0.0) 1.00 1.00 (0.0) 0.38 (0.0) 0.99 1.03 (0.0) 1.00 1.00 (0.0)
e5 gran. 0.81 (0.0) 0.99 2.20 (0.2) 0.02 (0.0) 0.00 4120.60 (9.3) 0.00 4096.36 (9.2)
gtr 0.74 (0.0) 0.82 2.56 (0.0) 0.03 (0.0) 0.00 4080.76 (9.3) 0.00 4065.74 (9.2)
gte 0.90 (0.0) 1.00 1.01 (0.0) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
stel. 0.78 (0.0) 1.00 1.00 (0.0) 0.38 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
@@
Table 2: In-distribution translations: vec2vecs trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 8192). The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses. Bold denotes best value.
@@
vec2vec translations mirror target geometry. Table˜2 shows that vec2vec generates embeddings with near-optimal assignment across model pairs, achieving cosine similarity scores up to 0.92, top-1 accuracies up to 100%, and ranks as low as 1. In same-backbone pairings (e.g., (gte, e5)), vec2vec’s top-1 accuracy and rank are comparable to both the naïve baseline and (surprisingly) the oracle-aided optimal assignment. Although the embeddings generated by vec2vec are significantly closer to the ground truth than the naïve baseline, in same-backbone pairings the embeddings are close enough to be compatible. In cross-backbone pairings, vec2vec is far superior on all metrics, while baseline methods perform similarly to random guessing.
@@
Table˜3 shows that this performance extends to out-of-distribution data. Our vec2vec translators were trained on NQ (drawn from Wikipedia), yet exhibit high cosine similarity, high accuracy, and low rank when evaluated on tweets (which are far more colloquial and use emojis) and medical records (which contain domain-specific jargon unlikely to appear in NQ). In Appendix˜B, we show that baseline methods fail on cross-backbone embedding pairs.
vec2vec OA Baseline
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ Top-1 $\uparrow$ Rank $\downarrow$
gran. clip 0.78 (0.0) 0.35 226.62 (3.2) 0.00 4096.32 (9.2)
gtr 0.73 (0.0) 0.13 711.23 (5.9) 0.00 4096.71 (9.2)
gte 0.62 (0.0) 0.00 3233.41 (9.8) 0.00 4096.32 (9.2)
stel. 0.77 (0.0) 0.31 286.69 (3.6) 0.00 4096.60 (9.2)
e5 0.64 (0.0) 0.01 2568.21 (9.4) 0.00 4096.78 (9.2)
clip gran. 0.74 (0.0) 0.72 4.46 (0.1) 0.00 4096.36 (9.2)
gtr 0.67 (0.0) 0.27 155.11 (2.1) 0.00 4096.80 (9.2)
gte 0.75 (0.0) 0.00 2678.90 (8.9) 0.00 4096.12 (9.2)
stel. 0.72 (0.0) 0.61 22.50 (0.5) 0.00 4096.47 (9.2)
e5 0.73 (0.0) 0.01 1692.28 (8.2) 0.00 4096.44 (9.2)
Table 4: Translations between unimodal and multimodal (CLIP) embeddings: vec2vecs trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 8192). Rank varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since the embedding dimensionalities are different, the naive baseline does not apply. Bold denotes best value.
Finally, Table˜4 shows that vec2vec can even translate to and from the space of CLIP, a multimodal embedding model which was trained in part on image data. While the translations are not as strong as in Table˜2, vec2vec consistently outperforms the optimal assignment baseline. These results show that promise of our method at adapting to new modalities: in particular, the embedding space of CLIP has been successfully connected to other modalities such as heatmaps, audio, and depth charts [11].
@@ 6 Using vec2vec translations to extract information
Zero-shot inversion. Inversion, i.e., reconstruction of text inputs, is more ambitious than attribute inference. vec2vec translations retain enough semantic information that off-the-shelf, zero-shot inversion methods like [59], developed for embeddings computed by standard encoders, extract information for as many as 80% of documents given only their translated embeddings, for some model pairs (Figure˜5). These inversions are imperfect and we leave development of specialized inverters for translated embeddings to future work. Nevertheless, as exemplified in Figure˜6, they still extract information such as individual and company names, dates, promotions, financial information, outages, and even lunch orders. In Appendix˜D, we show the prompt we use to measure extraction.

### v3

@@ Abstract
The ability to translate unknown embeddings into a different space while preserving their geometry has serious implications for security. An adversary with access to a database of only embedding vectors can extract sensitive information about underlying documents, sufficient for classification and attribute inference.
@@ 1 Introduction
The Platonic Representation Hypothesis [18] conjectures that all image models of sufficient size converge to the same latent representation. We propose a stronger, constructive version of this hypothesis for text models: the universal latent structure of text representations can be learned and, furthermore, harnessed to translate representations from one space to another without any paired data or encoders.
@@
vec2vec is the first method to successfully translate embeddings from the space of one model to another without paired data. [^1] Prior work has successfully translated word embeddings between languages, typically relying on overlapping vocabularies across languages. In contrast, we translate embeddings of entire sequences between model spaces. vec2vec translations achieve cosine similarity as high as $0.92$ to the ground-truth vectors in their target embedding spaces and perfect matching on over $8000$ shuffled embeddings (without access to the set of possible matches in advance).
To show that our translations preserve not only the relative geometry of embeddings but also the semantics of underlying inputs, we extract information from them using zero-shot attribute inference and inversion, without any knowledge of the model that produced the original embeddings. [^2] Our code is available on GitHub.
@@ 2 Problem formulation: unsupervised embedding translation
Figure 3: Unsupervised embedding translation. With access to only $u_{i}=M_{1}(d_{i})$, vec2vec seeks to generate a translation $F(u_{i})$ that is close in $M_{2}$’s embedding space to the ideal embedding $v_{i}=M_{2}(d_{i})$ without access to $d_{i}$, $v_{i}$, or $M_{1}$.
@@
Limitations of correspondence methods. There is significant prior research on the problem of matching or correspondence between sets of embedding vectors [1, 48, 8, 53]. These methods typically assume that the two (or more) sets of embeddings are generated by different encoders on the same or highly-overlapping inputs. In other words, for each unknown vector, there must already exist a set of candidate vectors in a different embedding. In practice, it is unrealistic to expect that such a database be available, so these methods are not directly applicable. Some matching methods, however, support translation between embedding spaces without overlapping inputs. Our experiments demonstrate that these methods struggle significantly, even when correspondence exists.
Our task is inherently more challenging than matching, because we do not assume access to encoder $M_{1}$, nor do we have additional representations of documents $d_{1},\ldots,d_{n}$} beyond their embeddings $u_{i}=M_{1}(d_{i})$. Therefore, we rely solely on unsupervised translation from $M_{1}$ to $M_{2}$. The effectiveness of such unsupervised translation approaches thus critically depends on identifying and leveraging shared geometric structures within the embedding spaces produced by $M_{1}$ and $M_{2}$.
@@ 4.1 Preliminaries
Training. Unless otherwise specified, each vec2vec is trained on two sets of embeddings generated from disjoint sets of $1$ million $64$-token sequences sampled from NQ (see Appendix˜D for experiments with fewer embeddings). Due to GAN instability [52], we select the best of multiple initializations and leave more robust training to future work.
@@ 4.2 Evaluating translation
Let $u_{i}=M_{1}(d_{i})$ and $v_{i}=M_{2}(d_{i})$ denote the source and target embeddings of the same input $d_{i}$. The goal of translation is to generate a vector that is as close to $v_{i}$ as possible. We say that $(u_{i},v_{j})$ are “aligned” by the translator $F$ if $v_{j}$ is the closest embedding to $F(u_{i})$: $j=\arg\min_{k}\cos\bigl(F(u_{i}),v_{k}\bigr).$ A perfect translator $F^{*}$ satisfies $i=\arg\min_{k}\cos\bigl(F^{*}(u_{i}),v_{k}\bigr)$ for all $i$.
@@
vec2vec is the first unsupervised embedding translator, thus there is no direct baseline. As our Naïve baseline, we simply use $F(x)=x$ to measure geometric similarity between embedding spaces. The second (pseudo)baseline is Oracle-aided optimal transport. It assumes that candidate targets are known and is thus strictly easier than vec2vec and the Naïve baseline. We solve optimal assignment, $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{n}\cos(u_{i},v_{\pi(i)})$, via either the Hungarian, Earth Mover’s Distance, Sinkhorn, or Gromov-Wasserstein algorithms, choosing the solver with the lowest rank for each experiment.
@@ 4.3 Evaluating information extraction
The second method is embedding inversion that recovers text inputs from embeddings. Since [42] requires a pre-trained inversion model for each embedding space, we use [66] instead to generate an approximation $d^{\prime}_{i}$ of $d_{i}$ from $F(M_{1}(d_{i}))$ in a zero-shot manner. We measure the extracted information using LLM judge accuracy: the fraction of translated embeddings for which GPT-4o determines that $d^{\prime}$ reveals information in $d$. See Appendix˜E for our prompt.
@@
vec2vec Naïve Baseline OT Baseline
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$
gra. gtr 0.80 (0.0) 0.99 1.19 (0.1) -0.03 (0.0) 0.00 4168.73 (9.2) 0.50 (0.0) 0.00 4094.22 (9.2)
gte 0.87 (0.0) 0.95 1.18 (0.0) 0.01 (0.0) 0.00 4088.58 (9.2) 0.85 (0.0) 0.00 4069.91 (9.2)
ste. 0.79 (0.0) 0.98 1.05 (0.0) 0.01 (0.0) 0.00 4208.26 (9.2) 0.45 (0.0) 0.00 4096.35 (9.2)
e5 0.85 (0.0) 0.98 1.11 (0.0) 0.02 (0.0) 0.00 4111.60 (9.2) 0.68 (0.0) 0.00 4096.17 (9.2)
gtr gra. 0.81 (0.0) 0.99 1.02 (0.0) -0.03 (0.0) 0.00 4169.76 (9.2) 0.50 (0.0) 0.00 4093.55 (9.2)
gte 0.87 (0.0) 0.93 2.31 (0.1) 0.04 (0.0) 0.00 4080.92 (9.2) 0.85 (0.0) 0.00 4079.92 (9.2)
ste. 0.80 (0.0) 0.99 1.03 (0.0) 0.00 (0.0) 0.00 4198.78 (9.2) 0.46 (0.0) 0.00 4093.85 (9.2)
e5 0.83 (0.0) 0.84 2.88 (0.2) 0.03 (0.0) 0.00 4082.84 (9.2) 0.83 (0.0) 0.00 4066.42 (9.2)
gte gra. 0.75 (0.0) 0.95 1.22 (0.0) 0.01 (0.0) 0.00 4079.81 (9.3) 0.69 (0.0) 0.00 4069.23 (9.2)
gtr 0.75 (0.0) 0.91 2.64 (0.1) 0.04 (0.0) 0.00 4084.15 (9.2) 0.70 (0.0) 0.00 4078.45 (9.2)
ste. 0.89 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 0.69 (0.0) 1.00 1.00 (0.0)
e5 0.87 (0.0) 0.99 5.19 (0.5) 0.68 (0.0) 1.00 1.00 (0.0) 0.83 (0.0) 1.00 1.00 (0.0)
ste. gra. 0.80 (0.0) 0.98 1.08 (0.0) 0.01 (0.0) 0.00 4209.08 (9.3) 0.48 (0.0) 0.00 4096.38 (9.2)
gtr 0.82 (0.0) 1.00 1.10 (0.0) 0.00 (0.0) 0.00 4192.31 (9.2) 0.50 (0.0) 0.00 4095.46 (9.2)
gte 0.92 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 0.86 (0.0) 1.00 1.00 (0.0)
e5 0.86 (0.0) 1.00 1.00 (0.0) 0.38 (0.0) 0.99 1.03 (0.0) 0.83 (0.0) 1.00 1.00 (0.0)
e5 gra. 0.81 (0.0) 0.99 2.20 (0.2) 0.02 (0.0) 0.00 4120.60 (9.3) 0.69 (0.0) 0.00 4096.12 (9.2)
gtr 0.74 (0.0) 0.82 2.56 (0.0) 0.03 (0.0) 0.00 4080.76 (9.3) 0.70 (0.0) 0.00 4065.74 (9.2)
gte 0.90 (0.0) 1.00 1.01 (0.0) 0.68 (0.0) 1.00 1.00 (0.0) 0.86 (0.0) 1.00 1.00 (0.0)
ste. 0.78 (0.0) 1.00 1.00 (0.0) 0.38 (0.0) 1.00 1.00 (0.0) 0.68 (0.0) 1.00 1.00 (0.0)
@@
Table 2: In-distribution translations: vec2vecs trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 8192). The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Standard errors are shown in parentheses. Bold denotes best value. Symbols denote the lowest-rank solver for specific experiments: Sinkhorn and Gromov-Wasserstein.
@@
vec2vec translations mirror target geometry. Table˜2 shows that vec2vec generates embeddings with near-optimal assignment across model pairs, achieving cosine similarity scores up to 0.92, top-1 accuracies up to 100%, and ranks as low as 1. In same-backbone pairings (e.g., (gte, e5)), vec2vec’s top-1 accuracy and rank are comparable to both the naïve baseline and (surprisingly) the oracle-aided optimal transport. Although the embeddings generated by vec2vec are significantly closer to the ground truth than the naïve baseline, in same-backbone pairings the embeddings are close enough to be compatible. In cross-backbone pairings, vec2vec is far superior on all metrics, while baseline methods perform similarly to random guessing.
@@
Table˜3 shows that this performance extends to out-of-distribution data. Our vec2vec translators were trained on NQ (drawn from Wikipedia), yet exhibit high cosine similarity, high accuracy, and low rank when evaluated on tweets (which are far more colloquial and use emojis) and medical records (which contain domain-specific jargon unlikely to appear in NQ). In Appendix˜C, we show that baseline methods fail on cross-backbone embedding pairs.
vec2vec OT Baseline
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$
gra. clip 0.78 (0.0) 0.35 226.62 (3.2) 0.60 (0.0) 0.00 4096.00 (9.2)
gtr 0.73 (0.0) 0.13 711.23 (5.9) 0.59 (0.0) 0.00 4096.78 (9.2)
gte 0.62 (0.0) 0.00 3233.41 (9.8) 0.59 (0.0) 0.00 4096.16 (9.2)
ste. 0.77 (0.0) 0.31 286.69 (3.6) 0.59 (0.0) 0.00 4096.76 (9.2)
e5 0.64 (0.0) 0.01 2568.21 (9.4) 0.60 (0.0) 0.00 4096.41 (9.2)
clip gra. 0.74 (0.0) 0.72 4.46 (0.1) 0.48 (0.0) 0.00 4096.86 (9.2)
gtr 0.67 (0.0) 0.27 155.11 (2.1) 0.49 (0.0) 0.00 4096.35 (9.2)
gte 0.75 (0.0) 0.00 2678.90 (8.9) 0.72 (0.0) 0.00 4096.44 (9.2)
ste. 0.72 (0.0) 0.61 22.50 (0.5) 0.45 (0.0) 0.00 4096.27 (9.2)
e5 0.73 (0.0) 0.01 1692.28 (8.2) 0.68 (0.0) 0.00 4096.42 (9.2)
Table 4: Translations between unimodal and multimodal (CLIP) embeddings: vec2vecs trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 8192). Rank varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since the embedding dimensionalities are different, only the Gromov-Wasserstein OT baseline is run and the naive baseline does not apply. Bold denotes best value.
Finally, Table˜4 shows that vec2vec can even translate to and from the space of CLIP, a multimodal embedding model which was trained in part on image data. While the translations are not as strong as in Table˜2, vec2vec consistently outperforms the optimal transport baseline. These results show the promise of our method at adapting to new modalities: in particular, the embedding space of CLIP has been successfully connected to other modalities such as heatmaps, audio, and depth charts [12].
@@ 6 Using vec2vec translations to extract information
Zero-shot inversion. Inversion, i.e., reconstruction of text inputs, is more ambitious than attribute inference. vec2vec translations retain enough semantic information that off-the-shelf, zero-shot inversion methods like [66], developed for embeddings computed by standard encoders, extract information for as many as 80% of documents given only their translated embeddings, for some model pairs (Figure˜5). These inversions are imperfect and we leave development of specialized inverters for translated embeddings to future work. Nevertheless, as exemplified in Figure˜6, they still extract information such as individual and company names, dates, promotions, financial information, outages, and even lunch orders. In Appendix˜E, we show the prompt we use to measure extraction.